OpenStudy (anonymous):

what is the period of y=tan(3x)

OpenStudy (xapproachesinfinity):

The period of tanx is pi so what will be the period for tan3x?

OpenStudy (anonymous):

pi over 3? im not sure...

OpenStudy (anonymous):

3x

OpenStudy (anonymous):

3pi

OpenStudy (xapproachesinfinity):

well here as situation where that tan(x)=0 it happens for every multiple of pi

OpenStudy (anonymous):

....

OpenStudy (anonymous):

i dont get it

OpenStudy (xapproachesinfinity):

you already answered!

OpenStudy (jhannybean):

Try drawing a picture.

OpenStudy (anonymous):

oh wowww

OpenStudy (anonymous):

OH WAIT I GET IT

OpenStudy (anonymous):

wait idk lol...

OpenStudy (anonymous):

tangent is on its period? :O

OpenStudy (xapproachesinfinity):

I'm just giving more sense to your answer

OpenStudy (xapproachesinfinity):

No you know, just what you said first

OpenStudy (anonymous):

i just guessed

OpenStudy (anonymous):

pi over 3?

OpenStudy (anonymous):

i feel like its that

OpenStudy (xapproachesinfinity):

well that is why i give you some more evidence to make sure

OpenStudy (xapproachesinfinity):

yes!

OpenStudy (anonymous):

The correct answer is pi/3.

OpenStudy (anonymous):

how about when i do y=cot(11/3 pi)

OpenStudy (anonymous):

is it 3/11

OpenStudy (anonymous):

or is it negative?

OpenStudy (xapproachesinfinity):

you are looking for what?

OpenStudy (anonymous):

period

OpenStudy (anonymous):

It is pi/3 No negatives bud

OpenStudy (anonymous):

ohhhh but why? what happens to the 11

OpenStudy (xapproachesinfinity):

is it y=cot(11/3x)?

OpenStudy (anonymous):

yessss sorry

OpenStudy (xapproachesinfinity):

I don't know why you are looking for the period in cot(11pi/3)? if you are asking y=cot(11/3x) it is the same concept cot have a period of pi like tan

OpenStudy (xapproachesinfinity):

So yeah you are right in your first attempt 3/11pi is good

OpenStudy (anonymous):

it was an example question thats why i was just making sure. thank you brother

OpenStudy (xapproachesinfinity):

Welcome! it just about dividing the period all the time

OpenStudy (anonymous):

its so weird though haha

OpenStudy (xapproachesinfinity):

hehe why? make sense! draw it and you will see what is happening

OpenStudy (anonymous):

that's true, and i have one last question, how does it work for sec?

OpenStudy (anonymous):

like trying to find the period

OpenStudy (xapproachesinfinity):

sec have period of 2pi like cosine

OpenStudy (xapproachesinfinity):

every 2pi it take the same values that's what happens

OpenStudy (anonymous):

that confuses me... i have one last question that is y=sec(pi/9 x)

OpenStudy (anonymous):

but i know the period is 9

OpenStudy (anonymous):

isn't**

OpenStudy (xapproachesinfinity):

pi/9x=2pi

OpenStudy (anonymous):

???? thats so weird lol

OpenStudy (xapproachesinfinity):

lol why? 18 radians

OpenStudy (anonymous):

the period is 2pi?

OpenStudy (xapproachesinfinity):

no 18 radians

OpenStudy (anonymous):

ohhhhh

OpenStudy (xapproachesinfinity):

not a full circle not a half circle but 18 radians

OpenStudy (anonymous):

what is 14 radians

OpenStudy (anonymous):

would that be 2pi too

OpenStudy (xapproachesinfinity):

14 radians in degrees?

OpenStudy (xapproachesinfinity):

No it is not

OpenStudy (anonymous):

oh im only asking cause i had to find period of csc(7x) so i thought it would be 1/2

OpenStudy (xapproachesinfinity):

7x=2pi ====> x=2pi/7

OpenStudy (anonymous):

ohh

ganeshie8 (ganeshie8):

csc(7x) runs 7 times faster than the usual csc(x)

OpenStudy (anonymous):

thanks guys!!!

ganeshie8 (ganeshie8):

it helps to think of period as the "time taken" for completing one cycle

OpenStudy (xapproachesinfinity):

@ganeshie8 Exactly!

OpenStudy (xapproachesinfinity):

the bigger the amplitude "i think its called" the faster it is cos(At) A amplitude

jimthompson5910 (jim_thompson5910):

I'm a bit late to the party, but the period of y = A*tan(Bx-C) + D is T = pi/B this is for the first problem at the top.

ganeshie8 (ganeshie8):